# CodeFuse-CR-Bench: A Comprehensiveness-aware Benchmark for End-to-End Code Review Evaluation in Python Projects

### v1

@@ 4.2.2 Semantics Similarity
In order to evaluate the semantics performance of review comments, we use BLEU Papineni et al. (2002), a popular evaluation metric that is implemented in neural machine translation and conversation systems. BLEU is now widely used in code-related tasks, such as code comment generation Guo et al. (2023), document generation Hu et al. (2022b), review comment generation Li et al. (2022b) and so on. Its core idea is to compare the n-gram overlap between the text generated by the model and the ground truth. The greater the overlap, the higher the BLEU score. The specific calculation process of BLEU is given as follows:
@@ 5.2 Studied LLMs
We consider mainstream LLMs—encompassing both open-source and proprietary models—that have been widely adopted in recent software-engineering-related studies Fan et al. (2025); Batole et al. (2025); Wang et al. (2024); Yin et al. (2024). For open-source LLMs, we select DeepSeek-v3.1 Guo et al. (2025), Kimi-K2-0905-preview Team (2025) Qwen3-235B-A22B Yang et al. (2025). For closed-source LLMs, we choose the commonly used commercial models: Claude-Sonnet-4-20250514 Anthropic (2025), Gemini 2.5 Pro Comanici et al. (2025), GPT-4o Hurst et al. (2024), GPT-5 OpenAI (2025a). They have excellent performance in code-related tasks in previous benchmarks.
@@ 5.3 Studied Context Acquisition Strategies
As a repository-level CR benchmark, CodeFuse-CR-Bench can provide repository-level contextual information as part of the LLM input to help the LLM achieve comprehensive CR. Due to the large size of modern software repositories, it is not feasible to feed the entire repository to the model. It is necessary to retrieve the most relevant information. Referring to SWE-Bench Jimenez et al. (2024), we adopt two types of context acquisition strategies that might be beneficial for comprehensive CR: Retrieval-based Acquisition and Oracle-based context acquisition. In RQ3, we will validate the performance of these two strategies.
Retrieval-based Context Acquisition. Since the problem description and the code implementing the functionality may share overlapping vocabulary, we employ BM25 Robertson et al. (2004) to retrieve code files similar to the problem statement within the project. BM25 is a ranking function commonly used in information retrieval to measure the relevance between a document and a query. As a kind of sparse retrieval method, BM25 offers fast computation speed and low resource consumption, making it highly suitable for file retrieval scenarios in large-scale software projects. In RQ3, we will separately analyze the impact of using the top-1, top-3 and top-5 relevant code files retrieved via BM25 as context on the effectiveness of LLM-based CR.

### v2

@@ 4.2.2 Semantics Similarity
In order to evaluate the semantics performance of review comments, we use BLEU (Papineni et al., 2002), a popular evaluation metric that is implemented in neural machine translation and conversation systems. BLEU is now widely used in code-related tasks, such as code comment generation (Guo et al., 2023), document generation (Hu et al., 2022b), review comment generation (Li et al., 2022b) and so on. Its core idea is to compare the n-gram overlap between the text generated by the model and the ground truth. The greater the overlap, the higher the BLEU score. The specific calculation process of BLEU is given as follows:
@@ 5.2 Studied LLMs
We consider mainstream LLMs—encompassing both open-source and proprietary models—that have been widely adopted in recent software-engineering-related studies (Fan et al., 2025; Batole et al., 2025; Wang et al., 2024; Yin et al., 2024). For open-source LLMs, we select DeepSeek-v3.1 (Guo et al., 2025), Kimi-K2-0905-preview (Team, 2025) Qwen3-235B-A22B (Yang et al., 2025). For closed-source LLMs, we choose the commonly used commercial models: Claude-Sonnet-4-20250514 (Anthropic, 2025), Gemini 2.5 Pro (Comanici et al., 2025), GPT-4o (Hurst et al., 2024), GPT-5 (OpenAI, 2025a). They have excellent performance in code-related tasks in previous benchmarks.
@@ 5.3 Studied Context Acquisition Strategies
As a repository-level CR benchmark, CodeFuse-CR-Bench can provide repository-level contextual information as part of the LLM input to help the LLM achieve comprehensive CR. Due to the large size of modern software repositories, it is not feasible to feed the entire repository to the model. It is necessary to retrieve the most relevant information. Referring to SWE-Bench (Jimenez et al., 2024), we adopt two types of context acquisition strategies that might be beneficial for comprehensive CR: Retrieval-based Acquisition and Oracle-based context acquisition. In RQ3, we will validate the performance of these two strategies.
Retrieval-based Context Acquisition. Since the problem description and the code implementing the functionality may share overlapping vocabulary, we employ BM25 Stephen2004 to retrieve code files similar to the problem statement within the project. BM25 is a ranking function commonly used in information retrieval to measure the relevance between a document and a query. As a kind of sparse retrieval method, BM25 offers fast computation speed and low resource consumption, making it highly suitable for file retrieval scenarios in large-scale software projects. In RQ3, we will separately analyze the impact of using the top-1, top-3 and top-5 relevant code files retrieved via BM25 as context on the effectiveness of LLM-based CR.